### v1

# Basis Selection: Low-Rank Decomposition of Pretrained Large Language Models for Target Applications

Abstract
Large language models (LLMs) significantly enhance the performance of various applications, but they are computationally intensive and energy-demanding. This makes it challenging to deploy them on devices with limited resources, such as personal computers and mobile/wearable devices, and results in substantial inference costs in resource-rich environments like cloud servers. To extend the use of LLMs, we introduce a low-rank decomposition approach to effectively compress these models, tailored to the requirements of specific applications. We observe that LLMs pretrained on general datasets contain many redundant components not needed for particular applications. Our method focuses on identifying and removing these redundant parts, retaining only the necessary elements for the target applications. Specifically, we represent the weight matrices of LLMs as a linear combination of base components. We then prune the irrelevant bases and enhance the model with new bases beneficial for specific applications. Deep compression results on the Llama 2-7b and -13B models, conducted on target applications including mathematical reasoning and code generation, show that our method significantly reduces model size while maintaining comparable accuracy to state-of-the-art low-rank compression techniques.
1 Introduction
Large language models (LLMs) have significantly enhanced the performance of various applications in natural language processing, computer vision, and beyond. However, their large model sizes pose a bottleneck for many practical uses. The substantial computing resources required for LLM inference make it challenging to deploy them on devices with limited capabilities, such as personal computers and mobile/wearable devices. Moreover, even on hardware platforms with ample computing power, deploying LLMs consumes a significant amount of energy, raising concerns about sustainability. Therefore, it is essential to reduce the size of LLMs after pretraining to ease their computational demands and lower energy consumption.
Our approach leverages the relationship between pretrained models and target applications. LLMs are typically pretrained on extensive datasets covering diverse tasks, which often share common elements. This shared pretraining generates synergies that contribute to the exceptional performance of LLMs. However, the differences between these tasks also introduce redundancy in the models. For a specific target application, we can remove these redundant components and retain only the useful parts of the LLMs. This approach allows us to compress the models while maintaining their performance on target applications as much as possible. Many scenarios only require supporting a certain type of application, making this approach effective in reducing the computing resource requirement and lowering inference costs.
Figure 1: Key idea of Basis Selection: Identify the important bases for target applications and keep them during compression.
However, how do we identify the beneficial and redundant components of LLMs for a specific application? In this work, we address this problem through the lens of matrix factorization. Singular Value Decomposition (SVD) [9] factorizes a weight matrix $W$ into the product of three matrices $U$, $S$, and $V$, i.e., $W=USV^{\mathrm{T}}=\sum_{i}{s_{i}\mathbf{u}_{i}{\mathbf{v}_{i}^{\mathrm{T}}}}$, where $s_{i}$ are (positive) singular values, and $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ are column vectors of $U$ and $V$ with unit norms. A recent study by Ferrando and Voita [8] demonstrates that these column vectors $u_{i}$ and $v_{i}$ carry specific meanings. For example, for an attention head typically activated in code generation tasks in the LLaMA 2-7B model [17], factorizing its weight matrix $W_{O}$ and de-embedding the resultant column vectors $\textbf{u}_{i}$ reveals tokens such as userId, groupId, filename, and className [8]. Obviously, these column vectors are highly useful for code generation tasks, but they are less beneficial for other tasks, such as mathematical reasoning.
Inspired by this observation, we propose Basis Selection, a low-rank decomposition approach to effectively compress LLMs for target applications. Figure 1 illustrates the key idea of Basis Selection. We view each weight matrix in LLMs as a linear combination of bases $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$ with singular values $s_{i}$ as their weights. These bases are valuable representations stored in the pretrained model, learned from large pretraining datasets. For a target application, some bases are advantageous while many others are not. To select the bases beneficial for the target application, we propose retraining the singular values (i.e., the weights of the bases) while keeping the bases fixed, using the training set of the target application. After retraining, we prune the bases associated with small singular values, as they are less important for the target application, and retain only those with large singular values, which are most critical for the target application. This approach allows us to eliminate the redundant parts of the original LLMs and retain only the components essential for the target application. To handle the data distribution differences between the pretraining dataset and the target application, we also augment the model with new bases learned from the training set of the target application during the pruning process. This enables us to learn the new bases necessary for the target application that are absent in the pretrained model.
We evaluate Basis Selection on two models—Llama 2-7B and Llama 2-13B [17]—and two tasks—mathematical reasoning and code generation. We evaluate the pass@1 accuracy of the compressed models on GSM8K [6] and MATH [10] for the mathematical reasoning task and HumanEval [3] and MBPP [2] for the code generation task. Compared to state-of-the-art baselines, our approach demonstrates significantly better performance when the compression ratio  [^1] The compression ratio is defined as the ratio of the original model size to the compressed model size. exceeds 6 for mathematical reasoning and 4 for code generation. This also implies that, for deep compression, our approach reduces the model size more significantly while maintaining similar accuracy to baseline methods.
This paper makes the following critical contributions:
  • We analyze the relationship between pretrained models and target applications, highlighting the opportunity and underlying rationale for using low-rank decomposition to compress large language models while maintaining performance on target applications.
  • We propose Basis Selection, a low-rank decomposition approach to compress pretrained large language models for target applications. Basis Selection identifies the beneficial and redundant components of large language models by relearning the importance (i.e., singular values) of bases using the training set of the target application, and then selects bases based on their importance.
  • We evaluate Basis Selection across multiple tasks and models, demonstrating its superior performance in deep compression.
2 Related Work
Singular Value Decomposition (SVD) [9] has been applied to reduce the size of machine learning models. Prior research [18, 13, 7, 20, 15, 4, 1, 14] has developed various SVD algorithms to compress different components of models, such as DNNs, CNNs, and embedding layers for a range of applications including natural language processing, speech, and vision. The primary distinction between our work and these prior studies is that they do not relearn the importance of bases using the training data of target applications. Instead, they typically prune bases according to the singular values from the original or finetuned models. FWSVD [11] evaluates the importance of individual weights rather than bases during SVD. As highlighted in Sections 1 and 3, the bases hold significant physical meanings. Considering importance at this level of granularity results in improved performance. As shown in Section 4, our approach surpasses FWSVD in deep compression performance.
A recent study [16] applied SVD to large language models. Its focus is on determining the optimal rank for each layer, while our emphasis is on basis selection. The two methods are orthogonal but complementary and can be combined. Chen et al. [5] and Yu and Wu [19] suggest reconstructing bases based on feature mimicking. These approaches are orthogonal to ours—they concentrate on basis reconstruction, whereas we focus on basis selection. Their methods complement ours and can be integrated together to achieve enhanced compression results.
3 Basis Selection
In this section, we describe our proposed compression method, Basis Selection.
For a linear layer $\mathbf{y}=W\mathbf{x}+\mathbf{b}$, Singular Value Decomposition (SVD) factorizes its weight matrix $W\in\mathbb{R}^{n\times m}$ as the product of three matrices $U$, $S$, and $V$:
$W=USV^{\mathrm{T}}$ (1)
where $U=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $S=\mathrm{diag}\left(s_{1},\cdots,s_{r}\right)$, and $V=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$.
The values $\left\{s_{i}\in\mathbb{R},i=1,\cdots,r\right\}$ are positive singular values. [^2] We drop zero singular values and the corresponding columns of matrices $U$ and $V$. The vectors $\left\{\mathbf{u}_{i}\in\mathbb{R}^{n},i=1,\cdots,r\right\}$ and $\left\{\mathbf{v}_{i}\in\mathbb{R}^{m},i=1,\cdots,r\right\}$ are orthonormal, i.e., $\left\|\mathbf{u}_{i}\right\|=1$, $\left\|\mathbf{v}_{i}\right\|=1$, $\mathbf{u}_{i}\perp\mathbf{u}_{j}$, and $\mathbf{v}_{i}\perp\mathbf{v}_{j}$ if $i\neq j$.
Therefore, we can factorize matrix $W$ as the following series:
$W=\sum_{i=1}^{r}s_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$ (2)
Let matrix $W_{i}=\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$, then
$\left\|W_{i}\right\|=\sqrt{\mathrm{tr}\left(W_{i}^{\mathrm{T}}W_{i}\right)}=\sqrt{\mathrm{tr}\left(\mathbf{v}_{i}\mathbf{u}_{i}^{\mathrm{T}}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}\right)}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}\mathbf{v}_{i}^{\mathrm{T}})}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}^{\mathrm{T}}\mathbf{v}_{i})}=1$ (3)
$\left\langle W_{i},W_{j}\right\rangle=\mathrm{tr}\left(W_{i}^{\mathrm{T}}W_{j}\right)=\mathrm{tr}\left(\mathbf{v}_{i}\mathbf{u}_{i}^{\mathrm{T}}\mathbf{u}_{j}\mathbf{v}_{j}^{\mathrm{T}}\right)=0,\quad\text{if}\;i\neq j$ (4)
Therefore, $\left\{\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}},i=1,\cdots,r\right\}$ can be seen as a group of orthonormal bases in a subspace of $\mathbb{R}^{n\times m}$, and $\left\{s_{i},i=1,\cdots,r\right\}$ are their weights, making the weight matrix $W$ a linear combination of these bases.
This group of bases can be viewed as a series of filters that manipulate the input signal $\mathbf{x}$ to produce the output signal $\mathbf{y}$:
$\mathbf{y}=W\mathbf{x}+\mathbf{b}=\sum_{i=1}^{r}s_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}\mathbf{x}+\mathbf{b}=\sum_{i=1}^{r}s_{i}\left\langle\mathbf{x},\mathbf{v}_{i}\right\rangle\mathbf{u}_{i}+\mathbf{b}$ (5)
In other words, for each basis (i.e., filter) $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$, the similarity between the input signal $\mathbf{x}$ and the unit direction vector $\mathbf{v}_{i}$ is measured by their inner product. This inner product is then multiplied by the (positive) singular value $s_{i}$ to determine the weight for the unit direction vector $\mathbf{u}_{i}$. The output signal $\mathbf{y}$ is the weighted sum of $\mathbf{u}_{i}$. Figure 2 illustrates this interpretation of the role of bases from the perspective of signal processing.
Figure 2: An interpretation of the role of bases from the perspective of signal processing.
Large language models pretrained on diverse datasets have bases that carry various meanings. Ferrando and Voita [8] investigate these meanings by factorizing the $W_{O}$ matrix of certain attention layers in Llama 2-7B and de-embedding the $\mathbf{u}$ vector associated with bases having large singular values. Table 3 presents key results from their study. It is observed that some bases have very specific meanings, such as those related to code generation. For target applications other than that, these bases are likely irrelevant. When using such bases as filters for irrelevant target applications, either the filter is not activated (due to a small inner product $\left\langle\mathbf{x},\mathbf{v}_{i}\right\rangle$), or worse, the filter is activated and harmful information is mixed into the output, leading to performance degradation. This suggests that pruning these bases can compress the models with negligible performance loss or, in some cases, even improve performance in the target application.
Table 1: The meaning of bases in Llama 2-7B (data are sourced from [8]).
Decoder Layer, Attention Head, Basis ID (sorted by singular values in descending order) Top ten most probable tokens corresponding to the vector $\mathbf{u}$ of the basis
Layer 17, Head 25, $W_{O}$ matrix, Basis 1 userID, groupId, filename, _className, artifactId, _filename, _Хронология, className, <s>, _javax
Layer 19, Head 17, $W_{O}$ matrix, Basis 4 _return, return, _Return, _Returns, _returning, Return, _returns, {CJK*}UTF8gbsn, _returned, returns
In our approach, Basis Selection, we determine the importance of the bases from the pretrained model by retraining their singular values on the training set of the target application. The weight matrix $\widetilde{W}$ in Basis Selection is represented as:
$\widetilde{W}=\sum_{i=1}^{r}\tilde{{s}}_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}+\sum_{j=1}^{\tilde{r}}\tilde{\mathbf{u}}_{j}\tilde{\mathbf{v}}_{j}^{\mathrm{T}}$ (6)
In the first term, $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$ represents the original bases in the pretrained model. To assess their importance, we initialize their weights $\tilde{s}_{i}$ with their original singular values and then retrain these weights (while keeping the bases fixed) on the training set of the target application. The aim is that, after retraining, the bases important for the target application will have larger singular values, whereas those that are useless or detrimental will have zero or very small singular values. This allows us to identify and prune the less useful bases. Relearning the importance of bases for the target application distinguishes our approach from previous methods. Prior approaches either use the singular values in the original model [18, 13, 7, 20, 15, 4, 1, 14, 16] or assess the importance of weight parameters, other than the importance of the bases, to prune them [11]. They do not relearn the importance of the bases specifically for the target application. From a signal processing perspective, this first term allows us to adjust the weight for each filter, catering to the needs of the target application.
In the second term, $\tilde{\mathbf{u}}_{j}$ and $\tilde{\mathbf{v}}_{j}$ are learnable vectors included for two primary purposes. First, due to differences in data distribution between the pretrained dataset and the target application, some bases necessary for the target application might be absent in the pretrained dataset. We use these vectors to learn such bases. Second, during pruning, although each pruned basis may individually have minimal impact on the target application, their cumulative performance loss can be significant. These additional vectors help compensate for the performance loss caused by the pruned bases. This second term is similar to LoRA [12]. The number of learnable vectors $\tilde{r}$ is referred to as the additional dimension. From a signal processing perspective, this second term allows us to include additional, new filters to enhance the model performance on the target application.
Algorithm 1 outlines our approach. Basis Selection takes a pretrained or fine-tuned model as input and gradually prunes bases in the original model. After each pruning step, it finetunes the learnable parameters $\tilde{s}_{i}$, $\tilde{\mathbf{u}}_{j}$, and $\tilde{\mathbf{v}}_{j}$ to offset the performance loss. Ultimately, a new weight matrix $\widetilde{W}$ with a smaller rank $r^{\prime}$ is learned. We perform a standard SVD on it, representing it as the product of matrices $U^{\prime}$, $S^{\prime}$, and $V^{\prime\mathrm{T}}$. We then replace the original layer with two new layers: $S^{\prime}V^{\prime\mathrm{T}}$ becomes the weight matrix of the first new layer, and $U^{\prime}$ becomes the weight matrix of the second new layer. This reduces the number of parameters from $nm$ to $(n+m)r^{\prime}$. The new model is subsequently further finetuned to enhance its performance on the target application.
Input: Pretrained or Finetuning Model $M$
Output: Compressed Model $M^{\prime}$
Data: KeepRatio, PruningTimes, KeepingEpoch, PruningEpoch, PostFineTuningEpoch
1
2IterationsPerPruning = round(NumIterationsPerEpoch * PruningEpoch / PruningTimes);
3 KeepRatioPerPruning = KeepRatio${}^{(1/\text{PruningTimes})}$;
4
5Convert the weight matrix of each layer in $M$ into the form of equation (6);
6 for $i=1$ to KeepingEpoch do
7       Tune the learnable parameters in equation (6);
8
9 end for
10
11for $i=1$ to PruningEpoch do
12       Tune the learnable parameters;
13       if IterationID is a multiple of IterationsPerPruning then
14             for each linear layer do
15                   Prune bases with smaller singular values $\tilde{s}_{i}$ such that after pruning, the sum of the singular values of the remaining bases is KeepRatioPerPruning of the sum before pruning;
16
17             end for
18
19       end if
20
21 end for
22
23for each layer do
24       Compute the low rank matrix $\widetilde{W}$ based on equation (6);
25       $[U^{\prime},S^{\prime},V^{\prime}]=\text{SVD}(\widetilde{W})$;
26       Use two linear layers to substitute for the original layer;
27       The first layer’s weight matrix is $S^{\prime}V^{\prime\textrm{T}}$;
28       The second layer’s weight matrix is $U^{\prime}$;
29
30 end for
31
32for $i=1$ to PostFineTuningEpoch do
33       FineTune the new model $M^{\prime}$;
34
35 end for
Algorithm 1 Basis Selection Algorithm
4 Experiments
(a) GSM8K
(b) MATH
Figure 3: Pass@1 accuracy and model size of Llama 2-7B compressed by various algorithms for the mathematical reasoning task.
(a) GSM8k
(b) MATH
Figure 4: Pass@1 accuracy and model size of Llama 2-13B compressed by various algorithms for the mathematical reasoning task.
4.1 Evaluation Methodology
The performance of low-rank compression algorithms is evaluated on two tasks: mathematical reasoning and code generation. For each task, Llama 2-7B and Llama 2-13B models [17] are first finetuned on a training dataset and then compressed using a compression algorithm. The compressed models are further finetuned before evaluation.
For the mathematical reasoning task, we utilize two evaluation datasets: GSM8K [6] and Hendrycks’ MATH [10]. The GSM8K dataset comprises verbally described mathematical questions, containing 1,319 samples used for evaluation. The Hendrycks’ MATH dataset covers more complex topics such as linear algebra and geometry, consisting of 5,000 question-answer pairs used for evaluation. Due to its complexity, the Hendrycks’ MATH dataset necessitates more sophisticated computations and reasoning, resulting in lower accuracy compared to GSM8K.
For the code generation task, we use two evaluation datasets: MBPP [2] and HumanEval [3]. Both datasets evaluate the models’ ability to generate Python code. MBPP comprises 500 code generation questions, while HumanEval includes 164 code generation questions.
We compare our proposed Basis Selection with state-of-the-art low-rank compression algorithms, specifically SVD and FWSVD. SVD is widely used in previous model compression studies [16, 1, 14, 18, 13, 7, 20, 15]. FWSVD [11] enhances SVD by evaluating the importance of model weight parameters.
Our work focuses exclusively on low-rank compression. Comparing our approach with model compression algorithms from other areas, such as quantization, sparsity & pruning, and knowledge distillation, is beyond the scope of this study. Ultimately, compression algorithms from different areas can be combined to achieve better compression results.
(a) HumanEval
(b) MBPP
Figure 5: Pass@1 accuracy and model size of Llama 2-7B compressed by various algorithms for the code generation task.
(a) HumanEval
(b) MBPP
Figure 6: Pass@1 accuracy and model size of Llama 2-13B compressed by various algorithms for the code generation task.
(a) GSM8k
(b) MATH
Figure 7: Ablation study: Effect of varying the additional dimension of Basis Selection on compressing Llama 2-7B for the mathematical reasoning task.
(a) GSM8k
(b) MATH
Figure 8: Ablation study: Effect of varying the pruning times of Basis Selection on compressing Llama 2-7B for the mathematical reasoning task.
4.2 Results on Mathematical Reasoning
Figures 3(a) and (b) depict the performance of various compression algorithms on the Llama 2-7B model for the mathematical reasoning task. We evaluate the models’ accuracy (Pass@1) at different compression ratios (original model size vs. compressed model size). For low compression ratios (below 6), all methods achieve similar accuracy. However, our Basis Selection method significantly outperforms SVD and FWSVD at higher compression ratios. For instance, at a 7x compression ratio, Basis Selection achieves around 43% and 10% accuracy on GSM8K and MATH datasets, respectively, while FWSVD drops below 2% accuracy on both datasets and SVD reaches only 27% and 5% accuracy on GSM8K and MATH, respectively. This highlights the effectiveness of Basis Selection for deep compression, especially when aiming for aggressive model size reduction.
Similar trends emerge for the larger Llama 2-13B model in Figures 4(a) and (b). Once again, Basis Selection significantly outperforms SVD and FWSVD at compression ratios exceeding 6. At a 7x compression ratio, Basis Selection achieves 47% and 11% accuracy on GSM8K and MATH datasets, respectively, demonstrating its advantage. This is in stark contrast to SVD’s performance (23% and 5% accuracy on GSM8K and MATH) and FWSVD’s near-complete accuracy drop (around 2% on both datasets). These results solidify Basis Selection’s effectiveness for deep compression across different model sizes.
4.3 Results on Code Generation
Similar results extend to code generation tasks (Figures 5 and  6). For both Llama 2-7B and Llama 2-13B models, all methods perform comparably at lower compression ratios (below 4). However, Basis Selection exhibits clear superiority at higher compression ratios. On Llama 2-7B at a 6x compression ratio, Basis Selection achieves 12% and 8% accuracy on HumanEval and MBPP datasets, respectively, significantly outperforming SVD (5% and 2%) and FWSVD (6% and 2%). Similar trends hold for Llama 2-13B. These findings further solidify Basis Selection’s effectiveness for deep compression across diverse tasks and model sizes.
4.4 Ablation Study
To analyze the impact of key parameters of Basis Selection, we conducted ablation studies on the additional dimension (denoted by $\tilde{r}$ in equation (6)) and pruning times. The additional dimension compensates for information loss during pruning, especially for deep compression. Figure 7 compares the performance of compressing Llama 2-7B for the mathematical reasoning task with and without an additional dimension of 32. As expected, incorporating an additional dimension improves accuracy at higher compression ratios (above 6). Similarly, pruning the model lightly but multiple times allows it to gradually adapt to the reduction in parameters after each pruning. This is particularly beneficial for achieving extreme compression ratios. Figure 8 demonstrates this effect by comparing the performance of pruning Llama 2-7B 100 times vs. 2 times on the same task. Here, pruning 100 times leads to better accuracy at compression ratios exceeding 10.
5 Conclusion
The significant size of large language models leads to high inference costs and demands substantial computing resources. To mitigate these issues, we focus on compressing large language models to meet the specific requirements of target applications. Our approach involves examining these models through the lens of matrix factorization. By viewing the weight matrix of large language models as a linear combination of a group of bases, we have identified that pretrained models often contain many redundant bases that are less useful for target applications. To address this, we propose Basis Selection, a compression algorithm that evaluates the importance of each base for target applications and prunes those that are less significant. Experimental results demonstrate that Basis Selection significantly outperforms state-of-the-art low-rank compression algorithms in achieving deep compression. We believe Basis Selection will greatly reduce the inference cost of large language models, making them more accessible and practical for a wider range of applications. This advancement has the potential to democratize the use of large language models, facilitating their adoption and integration across diverse fields and industries.